please thanks
2. 124 points Solve the following higher order homogeneous linear differential equatil y®+8y® +16y =0 1970 h quota 2y – 3y
Answer #1

- so auxiliary equation us m²-3 m²-8m-3=0 12m2-5m-3 )(m +i)=0 (2m+1)(m-3) (m+1=0 m= -1,-1,3 y=c, eat set Gezo 40+84.97164 =

go +246_344)_ _4 +44-o auxiliary equation as motams_3mt amet4m²=0 m² [m*+2m² - 3m² - 4m +4]=0 mo [ möcm +27 – 3m (m+2)+2(m+

I have solved the problem if you have any problem with the solution then please tell me on the comment box